# Optimal estimation for regression discontinuity design

### v1

@@ 3.1 Theoretical Comparisons
We compare the maximum MSE of the proposed estimator with that of the Gaussian estimator in the setting where the true model is the binary-outcome one considered in Section 2. In implementing the Gaussian estimator, the variance must be specified. In the following, we focus on the Gaussian estimator with $\sigma_{1}^{2}=\cdots=\sigma_{n}^{2}=1/4$ because the variance of a binary variable is less than or equal to $1/4$. Define
@@
Lemma 3.1 provides lower and upper bounds on the ratio of the maximum MSEs. Because $\hat{\bm{w}}$ minimizes $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\bm{w},\bm{\theta})$ over $\mathcal{W}$, the lower bound is trivial. In the proof of Lemma 3.1, we derive the upper bound by using an upper bound on the numerator and a lower bound on the denominator.
@@
To show the asymptotic result, we consider a uni-variate running variable $R_{i}$ and we assume that the running variable is bounded and the empirical distribution of $\|R_{i}\|$ is bounded above and below by linear functions. [^5] The convergence holds under a weaker condition which may be plausible for a multi-variate running variable. See Remark 3.1 for a discussion about the general case.
@@
    There exist $c_{1}>c_{0}>0$ such that, for any sufficiently large $n\in\mathbb{N}$, $c_{0}x-n^{-1/3}\leq F_{n}(x)\leq c_{1}x+n^{-1/3}$ for all $x\in[0,1]$, where $F_{n}(\cdot)$ is the empirical distribution of $\|R_{i}\|$ when the sample size is $n$, that is,
@@
Figure 3.2 illustrates Assumption 3.1 (ii). For example, when $R_{i}=i/n$ for all $i=1,\ldots,n$, this assumption is satisfied for $0<c_{0}<1<c_{1}$. This Assumption 3.1 (ii) requires that the empirical distribution $F_{n}(x)$ is bounded by a pair of linear functions.
Figure 3.2: The blue solid line denotes $y=F_{n}(x)$. The blue dotted lines denote functions $y=c_{1}x$ and $y=c_{0}x-\frac{1}{n}$.
@@
The convergence of $\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})$ holds under weaker restriction than Assumption 3.1. Specifically, the convergence holds for a multi-dimensional $R_{i}$. For example, suppose that for any $\epsilon>0$, the sample size satisfying $\|R_{i}\|\leq\epsilon$ goes to infinity as $n\to\infty$. That is, letting $N(\epsilon)\equiv\max\{i\in\{1,\ldots,n\}:\|R_{i}\|\leq\epsilon\}$, then $N(\epsilon)\to\infty$ holds for all $\epsilon>0$. This is weaker than Assumption 3.1 (ii) and plausible in a multi-dimentional case as well. In this case, for any $\epsilon>0$ we obtain
@@
where the first inequality is obtained by setting $\bm{w}=\left(\underbrace{\frac{1}{N(\epsilon)},\ldots,\frac{1}{N(\epsilon)}}_{N(\epsilon)},0,\ldots,0\right)^{\prime}$. Hence, $\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})\rightarrow 0$ as $\epsilon$ can be arbitrarily small.
@@
The shrinkage factor $\hat{u}=\sum_{i=1}^{n}\hat{w}_{i}$ converges to one under mild conditions. Consequently, the upper bound of Lemma 3.1 converges to $2$. To see this, we use the following relationship between $\hat{u}$ and $\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})$, which is the minimax MSE in the Gaussian model. In the proof of Theorem 3.1, we show that $\frac{1}{4}(1-\hat{u})^{2}\leq\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$. Because $\mathrm{MSE}(\bm{w},\bm{\theta})\leq\mathrm{MSE}_{g}(\bm{w},\bm{\theta})$ and $\Theta\subset\Theta_{g}$, we have
$\frac{1}{4}(1-\hat{u})^{2}\ \leq\ \max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})\ \leq\ \max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})\ \leq\ \max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta}).$ (3.3)
Hence, if $\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})$ converges to zero, the shrinkage factor $\hat{u}$ converges to one. From the discussion in Remark 3.1, we have $\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})\to 0$, and hence $\hat{u}\rightarrow 1$.
@@ 3.2 Numerical Comparisons
While the efficiency gain from our estimator relative to the Gaussian estimator can be small in large samples, their behaviors are quite different in finite samples. We demonstrate the finite-sample comparisons of our estimator with the Gaussian estimator in numerical analyses. Figures 3.3 and 3.4 plot weights $w_{1},\ldots,w_{n}$ for samples of observations whose values of the running variable are equally spaced between 0 and $1$. Figure 3.3 plots the weights of our estimator (rdbinary) and the Gaussian estimator (gauss) for the sample size of $50$ and four values of the Lipschitz constant. Figure 3.4 shows the plots for the sample size of $500$. The weights of the Gaussian estimator are computed under the assumption that the variance is homoskedastic and $1/4$ for the whole units as in Section 3.1. For the small sample size of $50$, our estimator exhibits moderate size of shrinkage whereas the Gaussian estimator has no shrinkage. For $C>0$, the weights of the Gaussian estimator are of a triangular shape, while the weights of our estimator have mild non-linearity. Also, the Gaussian weights have thicker tails than ours. These differences in shape arise from the fact that the Gaussian estimator is constructed under homoskedasticity and maximum possible variance of $1/4$, while ours optimizes the weights under potential heteroskedasticity.
Figure 3.3: Comparison of estimated weights for equally spaced grids (n = 50)
@@
On the other hand, the two estimators appear almost equivalent for a large enough sample size of $500$. The shape of our estimator remains sharper than the Gaussian estimator for $C=1$, but the differences between the two weights are negligible compared to the case with the small sample size of $50$.
Figure 3.4: Comparison of estimated weights for equally spaced grids (n = 500)
@@
Further distinct differences are in the maximal root MSEs in small samples. Figure 3.5 demonstrates the ratio of the maximum root MSE of the Gaussian estimator with $\sigma_{i}^{2}=1/4$ to that of our estimator, calculated in the binary-outcome model. For a small sample size of $50$, the Gaussian estimator has $5\%$ to $20\%$ larger root MSEs than our estimator. Hence, our estimator gains substantial improvements relative to the Gaussian estimator in small samples.
@@
Nevertheless, the ratios shrink as the sample size becomes larger and the gaps shrink below $5\%$ for $N=500$. This property is consistent with the theoretical result that the ratio of the worst-case MSEs converges to 1 as the sample size increases. In summary, our estimator is substantially different from and superior to the Gaussian estimator in finite samples, while the two estimators behave similarly in large samples.
Figure 3.5: Maximum root MSE ratio of Gaussian to rdbinary
4 Uniformly valid finite sample inference
@@
In this section, we return to the original setup introduced in Section 2.1, where we observe both the treated sample $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$ and the untreated sample $\{Y_{i,-},R_{i,-}\}_{i=1}^{n_{-}}$. We propose an inference procedure with respect to $\tau\equiv f(1,R_{0})-f(0,R_{0})$ based on a given linear shrinkage estimator. Let $p_{i,+}\equiv f(1,R_{i,+})$, $p_{i,-}\equiv f(0,R_{i,-})$, and $R_{0,+}=R_{0,-}=0$ so that $Y_{i,+}$ and $Y_{i,-}$ follow Bernoulli distribution with parameters $p_{i,+}$ and $p_{i,-}$, respectively. Similar to the previous sections, we assume that $p_{i,+}$ and $p_{i,-}$ satisfy $\bm{p}_{+}\equiv(p_{0,+},p_{1,+},\ldots,p_{n_{+},+})^{\prime}\in\mathcal{P}_{+}$ and $\bm{p}_{-}\equiv(p_{0,-},p_{1,-},\ldots,p_{n_{-},-})^{\prime}\in\mathcal{P}_{-}$, where
@@
Our inference procedure is valid for any linear estimator with nonnegative weights (even if $\sum_{i=1}^{n_{+}}w_{i,+}>1$ or $\sum_{i=1}^{n_{-}}w_{i,-}>1$) when the outcome is binary. Hence, we can conduct an inference using the linear shrinkage estimator proposed in the previous sections. Nevertheless, the following argument does not apply for general bounded outcomes. In Appendix C, we consider an inference procedure for general bounded outcomes.
@@ 4.1 One-sided test
We provide confidence intervals that are valid in finite samples by inverting tests that are valid in finite samples uniformly over the Lipschitz class. We begin our analysis from a one-sided test. Using the uniformly valid one-sided test, we construct a uniformly valid two-sided test and confidence interval.
@@
where $\gamma$ is a critical value. The critical value $\gamma$ must satisfy $P_{\bm{p}}(\hat{\tau}-\tau_{0}>\gamma)\leq\alpha$ for any parameter $\bm{p}\equiv(\bm{p}_{+}^{\prime},\bm{p}_{-}^{\prime})^{\prime}\in\mathcal{P}_{\ast}\equiv\mathcal{P}_{+}\times\mathcal{P}_{-}$ satisfying $H_{0}$. Hence, we need to choose the critical value $\gamma^{\ast}(\tau_{0})$ satisfying
@@
where $\mathcal{P}(\tau_{0})\equiv\left\{\bm{p}\in\mathcal{P}_{\ast}:p_{0,+}-p_{0,-}=\tau_{0}\right\}$. This critical value $\gamma^{*}(\tau_{0})$ provides a uniformly valid one-sided test in finite samples.
@@
Theorem 4.1 is obtained by using first-order stochastic dominance. Suppose that $(Y_{1},\ldots,Y_{n})^{\prime}\in\{0,1\}^{n}$ and $(\tilde{Y}_{1},\ldots,\tilde{Y}_{n})^{\prime}\in\{0,1\}^{n}$ follow $n$-dimensional independent Bernoulli distributions with parameters $\bm{p}\in\mathbb{R}^{n}$ and $\tilde{\bm{p}}\in\mathbb{R}^{n}$, respectively, and each element of $\bm{p}$ is larger than or equal to that of $\tilde{\bm{p}}$. Then, if $w_{i}$ is nonnegative for all $i$, $\sum_{i=1}^{n}w_{i}Y_{i}$ has first-order stochastic dominance over $\sum_{i=1}^{n}w_{i}\tilde{Y}_{i}$. Hence, if we fix $p_{0,+}$ and $p_{0,-}$, then $P_{\bm{p}}(\hat{\tau}>\gamma)$ is maximized at $\bm{p}=\left(\tilde{\bm{p}}_{+}(p_{0,+})^{\prime},\tilde{\bm{p}}_{-}(p_{0,-})^{\prime}\right)^{\prime}$, namely, (4.2) holds.
From Theorem 4.1, we can obtain the critical value $\gamma^{\ast}(\tau_{0})$ satisfying (4.1) by using the following algorithm:
@@
Because the critical value $\gamma^{\ast}(\tau_{0})$ depends on the hypothetical value $\tau_{0}$, we need to calculate the critical value for each hypothetical value. We can show that the critical value $\gamma^{\ast}(\tau_{0})$ is increasing in the hypothetical value $\tau_{0}$. Suppose that $1\leq\tau_{0}\leq\tilde{\tau}_{0}\leq 1$ and $p_{0,+}-p_{0,-}=\tau_{0}$. Then, there exist $\tilde{p}_{0,+}$ and $\tilde{p}_{0,-}$ such that $\tilde{p}_{0,-}\leq p_{0,-}$, $\tilde{p}_{0,+}\geq p_{0,+}$, and $\tilde{p}_{0,+}-\tilde{p}_{0,+}=\tilde{\tau}_{0}$. From the argument similar to the proof of Theorem 4.1, we obtain
@@
This result implies that $\gamma^{\ast}(\tau_{0})$ is increasing in $\tau_{0}$. Hence, if the null hypothesis $H_{0}:\tau=\tilde{\tau}_{0}$ is rejected, then the null hypothesis $H_{0}:\tau=\tau_{0}$ must be rejected for any $\tau_{0}<\tilde{\tau}_{0}$.

### v2

@@ 3.1 Theoretical comparisons
We compare the maximum MSE of the proposed estimator with that of the Gaussian estimator in a setting in which the true model is the binary-outcome model described in Section 2. The Gaussian estimator requires the specification of variance. In the following, we focus on the Gaussian estimator with $\sigma_{1}^{2}=\cdots=\sigma_{n}^{2}=1/4$ because the variance of a binary variable is less than or equal to $1/4$. Define
@@
Lemma 3.1 provides lower and upper bounds of the ratio of the maximum MSEs. As $\hat{\bm{w}}$ minimizes $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\bm{w},\bm{\theta})$ over $\mathcal{W}$, the lower bound is trivial. In the proof of Lemma 3.1, we derive the upper bound based on an upper bound on the numerator and a lower bound on the denominator.
@@
To establish the asymptotic result, we consider a univariate running variable $R_{i}$ and assume that it is bounded and that the empirical distribution of $\|R_{i}\|$ is bounded above and below by linear functions. [^6] The convergence holds under a weaker condition, which may be plausible for a multivariate running variable. See Remark 3.1 for a discussion of the general case.
@@
    There exist constants $c_{1}>c_{0}>0$ such that, for any sufficiently large $n\in\mathbb{N}$, $c_{0}x-n^{-1/3}\leq F_{n}(x)\leq c_{1}x+n^{-1/3}$ for all $x\in[0,1]$, where $F_{n}(\cdot)$ is the empirical distribution of $\|R_{i}\|$ when the sample size is $n$, that is,
@@
Figure 3.2 illustrates Assumption 3.1 (ii). For example, when $R_{i}=i/n$ for all $i=1,\ldots,n$, this assumption is satisfied for $0<c_{0}<1<c_{1}$. More generally, Assumption 3.1 (ii) requires the empirical distribution $F_{n}(x)$ to be bounded by a pair of linear functions.
Figure 3.2: The blue solid line denotes $y=F_{n}(x)$. The blue dotted lines denote functions $y=c_{1}x+n^{-1/3}$ and $y=c_{0}x-n^{-1/3}$.
@@
The convergence of $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$ and $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})$ to zero requires weaker conditions than Assumption 3.1. Specifically, the convergence may hold for multidimensional $R_{i}$. For example, suppose that for any $\epsilon>0$, the sample size satisfying $\|R_{i}\|\leq\epsilon$ goes to infinity as $n\to\infty$. That is, letting $N(\epsilon)\equiv\max\{i\in\{1,\ldots,n\}:\|R_{i}\|\leq\epsilon\}$, $N(\epsilon)\to\infty$ holds for all $\epsilon>0$. This condition is weaker than Assumption 3.1 (ii) and is plausible in a multidimensional case as well. To show the convergence of $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$ and $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})$ under this condition, we use the following relationship:
$\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})\ \leq\ \max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})\ \leq\ \max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta}),$
where the second inequality holds since $\mathrm{MSE}(\bm{w},\bm{\theta})\leq\mathrm{MSE}_{g}(\bm{w},\bm{\theta})$ and $\Theta\subset\Theta_{g}$. For any $\epsilon>0$, we obtain
@@
where the first inequality is obtained by setting $\bm{w}=\left(\underbrace{\frac{1}{N(\epsilon)},\ldots,\frac{1}{N(\epsilon)}}_{N(\epsilon)},0,\ldots,0\right)^{\prime}$ and the convergence follows from the assumption that $N(\epsilon)\to\infty$. Hence, $\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})\rightarrow 0$ since $\epsilon$ can be arbitrarily small. Therefore, $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$ and $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})$ also converge to zero.
@@
The shrinkage factor $\hat{u}=\sum_{i=1}^{n}\hat{w}_{i}$ converges to one under mild conditions. Consequently, the upper bound $2\hat{u}^{-2}$ of the ratio of the maximum MSEs given in Lemma 3.1 converges to $2$. To see that $\hat{u}$ converges to one, we use the following relationship: $\frac{1}{4}(1-\hat{u})^{2}\leq\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$, as shown in the proof of Theorem 3.1. From the discussion in Remark 3.1, under mild conditions, we have $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})\rightarrow 0$, which implies the shrinkage factor $\hat{u}$ converges to one.
@@ 3.2 Numerical comparisons
Although the efficiency gain from our estimator relative to the Gaussian estimator can be small in large samples, their behaviors are quite different in finite samples. We demonstrate the finite-sample comparisons of our estimator with the Gaussian estimator through numerical analyses. Figures 3.3 and 3.4 plot the weights $w_{1},\ldots,w_{n}$ for samples of observations whose running-variable values are equally spaced between 0 and $1$. Figure 3.3 plots the weights of our estimator (rdbinary) and the Gaussian estimator (gauss) for a sample size of $50$ and four Lipschitz constant values. Figure 3.4 shows the corresponding plots for a sample size of $500$. The weights of the Gaussian estimator are computed under the assumption that the variance is homoskedastic and $1/4$ for all units, as in Section 3.1. For a small sample size of $50$, our estimator exhibits moderate size of shrinkage, whereas the Gaussian estimator exhibits no shrinkage. For $C>0$, the weights of the Gaussian estimator are triangular, while our estimator’s weights display mild nonlinearity. Also, the Gaussian weights have thicker tails than ours. These differences in shape arise because the Gaussian estimator is constructed under homoskedasticity and the maximum possible variance of $1/4$, whereas our estimator optimizes the weights under potential heteroskedasticity.
Figure 3.3: Comparison of weights for equally spaced grids (N = 50)
Figure 3.4: Comparison of weights for equally spaced grids (N = 500)
@@
By contrast, the two estimators appear almost equivalent for a sufficiently large sample size of $500$. The shape of our estimator’s weights is still sharper than that of the Gaussian weights for $C=1$; however, the differences between the two weights are negligible compared to the case with a small sample size of $50$.
Figure 3.5: Ratio of maximum root MSE of Gaussian to that of rdbinary
@@
More distinct differences appear in the maximum root MSEs in small samples. Figure 3.5 reports the ratio of the maximum root MSE of the Gaussian estimator to that of our estimator, calculated in the binary-outcome model. For a small sample size of $50$, the Gaussian estimator has $5\%$ to $20\%$ larger root MSEs than our estimator. Hence, our estimator achieves substantial improvement over the Gaussian estimator in small samples.
@@
Nevertheless, these ratios shrink as the sample size increases and the gaps shrink below $5\%$ when $N=500$. This property is consistent with the theoretical result that the ratio of the worst-case MSEs converges to one as the sample size increases. In summary, our estimator is substantially different from and superior to the Gaussian estimator in finite samples, whereas the two estimators behave similarly in large samples.
4 Uniformly valid finite-sample inference
@@
In this section, we return to the original setup introduced in Section 2.1, where we observe both the treated sample $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$ and the untreated sample $\{Y_{i,-},R_{i,-}\}_{i=1}^{n_{-}}$. We propose an inference procedure with respect to $\tau\equiv f(1,R_{0})-f(0,R_{0})$ based on a given linear shrinkage estimator. Let $p_{i,+}\equiv f(1,R_{i,+})$, $p_{i,-}\equiv f(0,R_{i,-})$, and $R_{0,+}=R_{0,-}=0$, so that $Y_{i,+}$ and $Y_{i,-}$ follow Bernoulli distribution with parameters $p_{i,+}$ and $p_{i,-}$, respectively. Similar to the previous sections, we assume that $p_{i,+}$ and $p_{i,-}$ satisfy $\bm{p}_{+}\equiv(p_{0,+},p_{1,+},\ldots,p_{n_{+},+})^{\prime}\in\mathcal{P}_{+}$ and $\bm{p}_{-}\equiv(p_{0,-},p_{1,-},\ldots,p_{n_{-},-})^{\prime}\in\mathcal{P}_{-}$, where
@@
Our inference procedure is valid for any linear estimator with nonnegative weights (even if $\sum_{i=1}^{n_{+}}w_{i,+}>1$ or $\sum_{i=1}^{n_{-}}w_{i,-}>1$) when the outcome is binary. Hence, we can conduct an inference using the linear shrinkage estimator proposed in the previous sections. The following argument does not extend to general bounded outcomes. In Appendix C, we consider an inference procedure for general bounded outcomes based on Hoeffding’s inequality.
@@ 4.1 One-sided test
We provide finite-sample valid confidence intervals by inverting tests that are uniformly valid over the Lipschitz class. We begin our analysis with a one-sided test. Using a uniformly valid one-sided test, we then construct a uniformly valid two-sided test and confidence interval.
@@
where $\gamma$ is a critical value. The critical value $\gamma$ must satisfy $P_{\bm{p}}(\hat{\tau}-\tau_{0}>\gamma)\leq\alpha$ for any parameter $\bm{p}\equiv(\bm{p}_{+}^{\prime},\bm{p}_{-}^{\prime})^{\prime}\in\mathcal{P}_{\ast}\equiv\mathcal{P}_{+}\times\mathcal{P}_{-}$ satisfying $H_{0}$. Hence, we must choose the critical value $\gamma^{\ast}(\tau_{0})$ such that
@@
where $\mathcal{P}(\tau_{0})\equiv\left\{\bm{p}\in\mathcal{P}_{\ast}:p_{0,+}-p_{0,-}=\tau_{0}\right\}$. The critical value $\gamma^{*}(\tau_{0})$ provides a uniformly valid one-sided test for finite samples.
@@
Theorem 4.1 is obtained using first-order stochastic dominance. Suppose that $(Y_{1},\ldots,Y_{n})^{\prime}\in\{0,1\}^{n}$ and $(\tilde{Y}_{1},\ldots,\tilde{Y}_{n})^{\prime}\in\{0,1\}^{n}$ follow $n$-dimensional independent Bernoulli distributions with parameters $\bm{p}\in\mathbb{R}^{n}$ and $\tilde{\bm{p}}\in\mathbb{R}^{n}$, respectively, and each element of $\bm{p}$ is smaller than or equal to that of $\tilde{\bm{p}}$. Then, if $w_{i}$ is nonnegative for all $i$, $\sum_{i=1}^{n}w_{i}\tilde{Y}_{i}$ has first-order stochastic dominance over $\sum_{i=1}^{n}w_{i}Y_{i}$. Hence, if we fix $p_{0,+}$ and $p_{0,-}$, then $P_{\bm{p}}(\hat{\tau}>\gamma)$ is maximized at $\bm{p}=\left(\tilde{\bm{p}}_{+}(p_{0,+})^{\prime},\tilde{\bm{p}}_{-}(p_{0,-})^{\prime}\right)^{\prime}$, implying (4.2) holds.
From Theorem 4.1, we can obtain the critical value $\gamma^{\ast}(\tau_{0})$ satisfying (4.1) using the following algorithm:
@@
As the critical value $\gamma^{\ast}(\tau_{0})$ depends on the hypothesized value $\tau_{0}$, the critical value must be calculated for each hypothesized value. We can show that the critical value $\gamma^{\ast}(\tau_{0})$ increases with the hypothesized value $\tau_{0}$. Suppose that $1\leq\tau_{0}\leq\tilde{\tau}_{0}\leq 1$ and $p_{0,+}-p_{0,-}=\tau_{0}$. Then, there exist $\tilde{p}_{0,+}$ and $\tilde{p}_{0,-}$ such that $\tilde{p}_{0,-}\leq p_{0,-}$, $\tilde{p}_{0,+}\geq p_{0,+}$, and $\tilde{p}_{0,+}-\tilde{p}_{0,+}=\tilde{\tau}_{0}$. From an argument similar to that in the proof of Theorem 4.1, we obtain
@@
This result implies that $\gamma^{\ast}(\tau_{0})$ is increasing in $\tau_{0}$. Hence, if the null hypothesis $H_{0}:\tau=\tilde{\tau}_{0}$ is rejected, then the null hypothesis $H_{0}:\tau=\tau_{0}$ must also be rejected for any $\tau_{0}<\tilde{\tau}_{0}$.